# Self-consistent Hartree-Fock-Bogoliubov approach for bosons: self-eliminating divergence and pure pair condensate

### v1

@@ III Energy functional and coupled equations for homogeneous state
where $n_{0}=b^{*}_{0}b_{0}/V$ is the single-particle condensate density. Next, from equations (21), we find the quantities $\xi_{\mathbf{p}}$ and $\Delta_{\mathbf{p}}$ in terms of the distribution functions $f_{\mathbf{p}}$ and $g_{\mathbf{p}}$. Eliminating the latter (see (23)) from the corresponding expressions and equation (26) and performing algebraic manipulations, we arrive at the following set of coupled equations for $\xi_{\mathbf{p}}$, $\Delta_{\mathbf{p}}$ and $b_{0}$ Tolmachev (1969); Peletminskii and Peletminskii (2010); Peletminskii et al. (2013):
@@
IV Contact interaction
It is clear that the system of equations (28d) is difficult to solve, even using numerical methods, due to its transcendental and integral nature. Therefore, we should make a certain assumption that simplifies its analysis while still allowing us to obtain reasonable physical results. Specifically, we employ the fact that at low temperatures, low-energy collisions are characterized by the scattering length $a$. Therefore, we can replace the realistic interatomic potential with an effective contact pseudopotential that reproduces the correct scattering length Pitaevskii and Stringari (2016),
@@ IV Contact interaction
Below, we are interested in the repulsive contact potential $U>0$. Therefore, since $\Delta>0$ and $I>0$, equation (31b) implies that $\cos(\gamma-2\beta)=1$. In addition, it is straightforward to see that replacing $U(\mathbf{p})$ with the contact interaction $U$ leads to a divergent term $I$. This problem is commonly known and also arises in other approaches Pitaevskii and Stringari (2016); Proukakis and Jackson (2008); Andersen (2004). Several ways have been proposed to address this problem: replacing the upper limit of integration with a cut-off value, using a non-contact model potential, renormalizing the coupling constant, or simply neglecting the divergent term.
Imposing a cut-off at sufficiently large momentum $p_{c}$ or employing a model potential with a finite range $r_{0}$ to eliminate the divergence, are in fact, equivalent procedures. Metaphorically speaking, applying a cut-off to the integral is as different from using a model potential as emergency braking is from service braking. In other words, the Fourier transforms of model potentials lead to a slow decay of the integrand, which ensures a convergence of the corresponding integrals Bulakhov et al. (2018). Nevertheless, a simple numerical analysis reveals that the solution to the equations is sensitive to small variations in $p_{c}$ or $r_{0}$, indicating that we cannot rely on the above methods to remove the divergence.
@@
IV.1 Gapless spectrum and Popov approximation
Any self-consistent solution of the coupled equations (31d) leads to a gapful spectrum of excitations. Nevertheless, let us demonstrate that neglecting the divergent term $I$ in the system (31d) yields a gapless Bogoliubov spectrum that satisfies the Pines–Hugenholtz condition Hugenholtz and Pines (1959). It is worth noting that this approximation is artificial and does not correspond to a self-consistent solution of (31d). Therefore, by neglecting the diverging term, one obtains
@@ IV.1 Gapless spectrum and Popov approximation
IV.2 Solution with self-eliminating divergence
@@
This implies that all correlated particles are in the condensate. At absolute zero temperature ($T=0$), all particles in the gas become simultaneously paired and condensed, giving rise to a macroscopically (correlated and coherent) quantum state (see Huang (1987), section 2.3).
@@
From (43), it is evident that $P^{\textrm{PA}}>P^{\textrm{SSED}}$. Furthermore, we obtain the interesting result that $P^{\textrm{SSED}}=0$ precisely at $T=0$, as the divergence is naturally eliminated. To explain this outcome, we follow the reasoning provided by Huang in his discussion of the Van der Waals gas Huang (1987), which served as an inspiration for understanding the vanishing pressure. Since the system of equations (31d) was derived self-consistently, the divergent term $I$ can be interpreted as an effective attractive interaction between an individual pair of particles, as also indicated by the negative sign in equation (31b). The physical origin of this attraction lies in the cumulative repulsion from all other particles in the gas acting on the considered pair. Recalling that pressure is defined as force per unit area, we observe that in our SSED framework, repulsive and attractive forces are exactly balanced at every point. Furthermore, another argument supporting the interpretation of $I$ as an attraction comes from the very form of the Bogoliubov pressure (44). Within the Bogoliubov approximation, the contribution of $I$ must be neglected, resulting in a finite positive pressure that reflects the dominance of repulsive interactions in the gas.
@@ IV.3 Normal state
In the ideal gas limit ($U=0$), the relations (53) and (54) completely reproduce the well-known expressions. Therefore, the terms $2nU$ in $\xi^{\textrm{NS}}_{\mathbf{p}}$ and $n^{2}U$ in $P^{\textrm{NS}}$ can be interpreted as interaction-induced corrections to the chemical potential and pressure, respectively, consistent with the results reported in Bulakhov et al. (2021); Ota and Giorgini (2020); Spada et al. (2022).
V Temperature dependence of physical observables
@@
Fig. 1 shows the temperature dependence of the condensate fraction for an ideal Bose gas, the PA and the SSED obtained from the coupled equations of the HFB approach. The curves for the PA and SSED are obtained by numerically solving equation (31d), which ensures the conservation of the total particle number. A notable feature of the presented dependencies is the retrograde behavior of the condensate density, which decreases to zero at $T=T^{\textrm{IBG}}$ following a reversal point at certain temperature $T=T_{r}$. While this behavior is qualitatively similar across different theoretical approaches (PA and SSED), the reversal temperature $T_{r}$ exhibits quantitative dependence on both the approximation used and the specific gas parameters. In particular, the dependencies of $T_{r}$ and the condensate fraction at this temperature on the gas parameter are demonstrated in Fig. 2. In the framework of PA, we observe a non-monotonic behavior of the condensate fraction at $T_{r}$, as also reported in Vianello and Salasnich (2024). The nontrivial temperature dependence of the condensate fraction is clearly a consequence of the nonlinear structure of equation (31d).
Figure 1: (Color online) Condensate fraction vs temperature for total density $n=1.6\cdot 10^{14}$ cm-3 and scattering length $a=4.9$ nm. The curves (from left to right) refer to: ideal Bose gas (IBG), solution with self-eliminating divergence (SSED) and Popov approximation (PA). Magnified sections of the SSED and PA curves with retrograde behaviour are shown in the inset.
@@
Figure 3: Temperature dependencies of: (a) chemical potential; (b) entropy; (c) pressure; and (d) pressure relative to that of an ideal Bose gas — for the total density of degenerate state $n^{\textrm{deg}}=1.6\cdot 10^{14}$ cm-3, the scattering length $a=4.9$ nm, $\mu_{c}=\mu^{\textrm{NS}}(T=T^{\textrm{IBG}})=2nU$, $S_{c}=S^{\textrm{NS}}(T=T^{\textrm{IBG}})=5/2\zeta(5/2)/\zeta(3/2)Nk_{B}$ and $P_{c}=P^{\textrm{NS}}(T=T^{\textrm{IBG}})=n^{2}U+\zeta(5/2)/\zeta(3/2)nk_{B}T^{\textrm{IBG}}$. The solid curves refer to: ideal Bose gas (IBG), solution with self-eliminating divergence (SSED), Popov approximation (PA) and normal state (NS) — for total density $n=n^{\textrm{deg}}$, i.e. relative density rd$=n/n^{\textrm{deg}}=1$. The dashed curves refer to the normal state for rd$=0.96$. Magnified sections of the foregoing curves with retrograde behaviour are shown in the insets.
Remarkably, despite the nonlinearity, chemical potentials and pressures of the degenerate and normal states coincide at $T=T^{\textrm{IBG}}$. Thus, the formal conditions for a phase transition are satisfied (see Huang (1987)). However, due to the retrograde behavior, this does not apply to the temperature derivatives of the foregoing quantities. Although the equality of derivatives at the transition point is not a necessary condition for the phase transition, the very existence of the derivatives is crucial. Indeed, the temperature derivative does not exist on the left side of the transition point, neither in the degenerate nor in the normal state. Therefore, the point $T=T^{\textrm{IBG}}$ cannot be identified as a genuine phase transition. Instead, the entire temperature interval $T\in[T^{\textrm{IBG}},T_{r}]$ should be considered as a region of an instability.
@@
Since the total number of particles is conserved while the densities differ between phases, the phase transition is accompanied by discontinuities in volume and entropy, as shown in Fig. 3(b). As is well known, this behavior corresponds to a first order phase transition Huang (1987). In particular, the entropy discontinuity leads to a nonzero latent heat. In addition, the order parameter also exhibits a discontinuous jump, as shown in Fig. 1. Employing this fact, we can clarify the volume discontinuity from a physical standpoint. Since the condensate does not contribute to the pressure, and a transition occurs at constant pressure, the loss of the latter is compensated by a decrease in volume.
@@
As we can see in Fig. 4, there is no crucial evidence (for example, $C>0$) pointing on any instability of the SSED compared to the PA. Neither the entropy nor the specific heat capacity in the normal state have linear corrections in interaction Akhiezer and Peletminskii (1981); Bulakhov et al. (2021). Therefore, these quantities coincide with those of the ideal Bose gas.
@@
Its temperature dependence is presented in Fig. 5, where the significant qualitative differences between the two approximations become apparent. The PA exhibits a positive and generally well-behaved compressibility (with divergence only in the region of instability), as also reported in Ota and Giorgini (2020); Mordini et al. (2020); Spada et al. (2022). In contrast, for the SSED, the compressibility is negative almost everywhere, diverging to minus infinity as $T\to 0$. Thus, the equilibrium state corresponding to the SSED is unstable.
@@
To summarize, we observe that almost all thermodynamic quantities exhibit well-behaved properties within both the HFB approach and the PA. Furthermore, the latter, which accounts solely for the single-particle condensate, predicts the positive compressibility. However, since the HFB approach does not admit a solution of (27) corresponding to a pure single-particle condensate Poluéktov (2002); Peletminskii et al. (2013), and the pair condensate (SSED) is unstable, the only possible stable state can be a mixture of single-particle and pair condensates , as also discussed below.
VI Discussion of results
@@
We have employed the energy-functional formulation to derive the self-consistency equations of the Hartree-Fock-Bogoliubov (HFB) approach. In the limit of contact interactions, these equations exhibit ultraviolet divergence, which are typically addressed by regularizing the corresponding integrals or renormalizing the coupling constant. We have obtained a solution to the HFB self-consistency equations in which the divergence self-eliminates (the solution with self-eliminating divergence, SSED). This solution possesses several distinct features. Specifically, it predicts the existence of a condensate in the form of correlated pairs only, with zero depletion and pressure at absolute zero temperature. The vanishing pressure can be explained by treating the divergent term as an effective attraction between an individual pair of particles, balanced by cumulative repulsion exerted by all other particles in the gas on the given pair. We have analyzed the temperature behavior of key thermodynamic quantities in the normal state and in the degenerate states described both by the obtained SSED and Popov approximation (PA). This analysis allows us to conclude that, due to interactions, the transition temperature shifts to higher values compared to that for an ideal gas, with the HFB approach leading to a larger shift than the PA. This positive shift of the transition temperature has also been reported in previous mean-field studies Vianello and Salasnich (2024) and Monte Carlo simulations Prokof’ev et al. (2004); Spada et al. (2022). In both cases, we propose to treat the phase transition between the normal and degenerate state as a first-order one. Unfortunately, the isothermic compressibility for SSED appears to be negative, which signals its instability. Finally, we emphasize that the differences among all the considered approaches, including the ideal gas case, become more pronounced as the temperature approaches $T^{\textrm{IBG}}$, which facilitates experimental verification.
@@
Some additional comments should also be made on certain important aspects of our study. First, although model non-contact potentials are not explicitly considered in this paper, some conclusions can nevertheless be drawn regarding their influence on the obtained results. As noted in Section IV, while they allow for a straightforward removal of divergences, the results lack stability with respect to variations in the interaction range and the choice of model potential. Furthermore, even within framework of Bogoliubov’s quadratic approximation, all solutions involving local potentials are also unstable in the same sense Bulakhov et al. (2018). Within the HFB approach, as indicated by (34), the gap lies in the range $E_{0}\in(0,2n_{0}U]$, since PA is gapless and SSED corresponds to $E_{0}=2n_{0}U$. This allows a qualitative conclusion that for non-contact potentials, the expected results will be bounded between the curves obtained from the SSED and the PA.
@@
In connection with the Nambu-Goldstone (NG) theorem Nambu (1960); Goldstone (1961) and the Pines–Hugenholtz relation Hugenholtz and Pines (1959), it is commonly believed that the presence of a gap in the single-particle excitation spectrum is a drawback of the self-consistent HFB approach. Numerous attempts have been made to artificially remove the gap in order to satisfy this relation. In particular, the most well-known gapless PA neglects anomalous averages, and in some literature sources Spada et al. (2022), the HFB approach is even considered to be valid only in the linear order in interaction. However, as we previously pointed out, these do not represent self-consistent solutions to the coupled HFB equations. We report a unique self-consistent solution to the HFB equations for a contact interaction that resolves the divergence issue without introducing additional assumptions or artificial regularization procedures. At sufficiently high temperatures, this solution exhibits qualitative agreement with the PA, except for the isothermal compressibility.
@@
Therefore, the gathered evidence encourages us to reconsider the nature of the energy gap in the HFB approach. On the one hand, the presence of gapless (phonon) excitations is not disputed. On the other hand, the HFB approach, along with the PA artificially derived from it, represent an attempt to construct a quasiparticle description of an interacting system in terms of a single type of single-particle excitations. This holds true if the spectra of single-particle and collective excitations coincide Gavoret and Nozières (1964). However, it has been recently demonstrated that the poles of the single-particle and two-particle Green’s functions in a Bose system with condensate do not coincide Kita (2009, 2011). Therefore, when both types of excitations predominate in the system, the application of the HFB framework leads to a unified spectrum of single-particle excitations hybridized with collective (pair) excitations, inheriting a gap from the latter. The presence of a gap in the two-particle excitations, however, is not an artifact, as it can be associated with the energy required to break a pair.
@@
The NG theorem states that the excitation spectrum arising in a system with spontaneously broken symmetry must be gapless. This is equivalent to the statement that it should cost zero energy to excite the lowest mode (the Goldstone mode). While the direct application of the NG theorem to quasiparticles, which are not physical particles, remains debated Sütő and Szépfalusy (2008), we highlight the following observation. For the SSED, where $\Delta=0$, the second equation in (21) shows that the energy change associated with the emergence of a single correlation is zero. Therefore, when pair correlations dominate in the ground state, as described by the solution we have found, the lowest mode is collective, and its excitation costs no energy.
@@
Thus, a system of interacting particles in the presence of a condensate requires a revised description that reproduces a spectrum with at least two excitation branches: a phonon-like mode and a gapped mode. Similar conclusions have been drawn in Refs. Bogolubov and Bogolubov Jr (2010); Poluektov and Soroka (2023). It should be emphasized that pure single-particle or pair condensates do not exist – only their mixture does. The existence of a purely single-particle condensate, as assumed in the PA, is ruled out by the system of HFB equations (27) (see also Poluéktov (2002); Peletminskii et al. (2013)). As for a pure pair condensate described by the SSED, it is found to be unstable based on the compressibility analysis. This result holds for the contact interaction model. In the case of finite-range interaction potentials, the corresponding HFB solutions also exhibit a gap, with $E_{0}<2n_{0}U$, and the ground state likewise represents a mixture of the two types of condensates. Such a structure of the ground state is also relevant in the context of superfluid helium, where the phonon-roton spectrum emerges from the hybridization of excitations of different physical natures and the single-particle condensate fraction is about 10% or less Sosnick et al. (1989); Bogoyavlenskii et al. (1990); Moroni and Boninsegni (2004); Glyde et al. (2011).

### v2

@@ 3 Energy functional and coupled equations for homogeneous state
where $n_{0}=b^{*}_{0}b_{0}/V$ is the single-particle condensate density. Next, from equations (21), we find the quantities $\xi_{\mathbf{p}}$ and $\Delta_{\mathbf{p}}$ in terms of the distribution functions $f_{\mathbf{p}}$ and $g_{\mathbf{p}}$. Eliminating the latter (see (23)) from the corresponding expressions and equation (26) and performing algebraic manipulations, we arrive at the following set of coupled equations for $\xi_{\mathbf{p}}$, $\Delta_{\mathbf{p}}$ and $b_{0}$ Tolmachev [1969], Peletminskii and Peletminskii [2010], Peletminskii et al. [2013]:
@@
4 Contact interaction
It is clear that the system of equations (28d) is difficult to solve, even using numerical methods, due to its transcendental and integral nature. Therefore, we should make a certain assumption that simplifies its analysis while still allowing us to obtain reasonable physical results. Specifically, we employ the fact that at low temperatures, low-energy collisions are characterized by the scattering length $a$. Therefore, we can replace the realistic interatomic potential with an effective contact pseudopotential that reproduces the correct scattering length Pitaevskii and Stringari [2016],
@@ 4 Contact interaction
Below, we are interested in the repulsive contact potential $U>0$. Therefore, since $\Delta>0$ and $I>0$, equations (29) and (31b) imply that $\cos(\gamma-2\beta)=1$. In addition, it is straightforward to see that replacing $U(\mathbf{p})$ with the contact interaction $U$ leads to a divergent term $I$. This problem is commonly known and also arises in other approaches Pitaevskii and Stringari [2016], Proukakis and Jackson [2008], Andersen [2004]. Several ways have been proposed to address this problem: replacing the upper limit of integration with a cut-off value, using a non-contact model potential, renormalizing the coupling constant, or simply neglecting the divergent term.
Imposing a cut-off at sufficiently large momentum $p_{c}$ or employing a model potential with a finite range $r_{0}$ to eliminate the divergence, are in fact, equivalent procedures. Metaphorically speaking, applying a cut-off to the integral is as different from using a model potential as emergency braking is from service braking. In other words, the Fourier transforms of model potentials lead to a slow decay of the integrand, which ensures a convergence of the corresponding integrals Bulakhov et al. [2018]. Nevertheless, a simple numerical analysis reveals that the solution to the equations is sensitive to small variations in $p_{c}$ or $r_{0}$, indicating that we cannot rely on the above methods to remove the divergence.
@@
4.1 Gapless spectrum and Popov approximation
Any self-consistent solution of the coupled equations (31d) leads to a gapful spectrum of excitations. Nevertheless, let us demonstrate that neglecting the divergent term $I$ in the system (31d) yields a gapless Bogoliubov spectrum that satisfies the Pines–Hugenholtz condition Hugenholtz and Pines [1959]. It is worth noting that this approximation is artificial and does not correspond to a self-consistent solution of (31d). Therefore, by neglecting the diverging term, one obtains
@@ 4.1 Gapless spectrum and Popov approximation
4.2 Solution with self-eliminating divergence
@@
This implies that all correlated particles are in the condensate. At absolute zero temperature ($T=0$), all particles in the gas become simultaneously paired and condensed, giving rise to a macroscopically (correlated and coherent) quantum state (see Huang [1987], section 2.3).
@@
From (43), it is evident that $P^{\textrm{PA}}>P^{\textrm{SSED}}$. Furthermore, we obtain the interesting result that $P^{\textrm{SSED}}=0$ precisely at $T=0$, as the divergence is naturally eliminated. To explain this outcome, we follow the reasoning provided by Huang in his discussion of the Van der Waals gas Huang [1987], which served as an inspiration for understanding the vanishing pressure. Since the system of equations (31d) was derived self-consistently, the divergent term $I$ can be interpreted as an effective attractive interaction between an individual pair of particles, as also indicated by the negative sign in equation (31b). The physical origin of this attraction lies in the cumulative repulsion from all other particles in the gas acting on the considered pair. Recalling that pressure is defined as force per unit area, we observe that in our SSED framework, repulsive and attractive forces are exactly balanced at every point. Furthermore, another argument supporting the interpretation of $I$ as an attraction comes from the very form of the Bogoliubov pressure (44). Within the Bogoliubov approximation, the contribution of $I$ must be neglected, resulting in a finite positive pressure that reflects the dominance of repulsive interactions in the gas.
@@ 4.3 Normal state
In the ideal gas limit ($U=0$), the relations (53) and (54) completely reproduce the well-known expressions. Therefore, the terms $2nU$ in $\xi^{\textrm{NS}}_{\mathbf{p}}$ and $n^{2}U$ in $P^{\textrm{NS}}$ can be interpreted as interaction-induced corrections to the chemical potential and pressure, respectively, consistent with the results reported in Bulakhov et al. [2021], Ota and Giorgini [2020], Spada et al. [2022].
5 Temperature dependence of physical observables
@@
Fig. 1 shows the temperature dependence of the condensate fraction for an ideal Bose gas, the PA and the SSED obtained from the coupled equations of the HFB approach. For comparison, Fig. 1 also displays Monte Carlo results taken from Ref. Prokof’ev et al. [2004]. The curves for the PA and SSED are obtained by numerically solving equation (31d), which ensures the conservation of the total particle number. A notable feature of the presented dependencies is the retrograde behavior of the condensate density, which decreases to zero at $T=T^{\textrm{IBG}}$ following a reversal point at certain temperature $T=T_{r}$. While this behavior is qualitatively similar across different theoretical approaches (PA and SSED), the reversal temperature $T_{r}$ exhibits quantitative dependence on both the approximation used and the specific gas parameters. In particular, the dependencies of $T_{r}$ and the condensate fraction at this temperature on the gas parameter are demonstrated in Fig. 2. In the framework of PA, we observe a non-monotonic behavior of the condensate fraction on interaction strength, as also reported in Vianello and Salasnich [2024]. The nontrivial temperature dependence of the condensate fraction is clearly a consequence of the nonlinear structure of equation (31d).
Figure 1: (Color online) Condensate fraction vs temperature for total density $n=1.6\cdot 10^{14}$ cm-3 and scattering length $a=4.9$ nm. The curves (from left to right) refer to: ideal Bose gas (IBG), Popov approximation (PA) and solution with self-eliminating divergence (SSED). Magnified sections of the SSED and PA curves with retrograde behaviour are shown in the inset. The scatter plot (MC) illustrates the predictions of the universal relations, calculated using the Monte-Carlo method Prokof’ev et al. [2004].
@@
[…]
Remarkably, despite the nonlinearity, chemical potentials and pressures of the degenerate and normal states coincide at $T=T^{\textrm{IBG}}$. Thus, the formal conditions for a phase transition are satisfied (see Huang [1987]). However, due to the retrograde behavior, this does not apply to the temperature derivatives of the foregoing quantities. Although the equality of derivatives at the transition point is not a necessary condition for the phase transition, the very existence of the derivatives is crucial. Indeed, the temperature derivative does not exist on the left side of the transition point, neither in the degenerate nor in the normal state. Therefore, the point $T=T^{\textrm{IBG}}$ cannot be identified as a genuine phase transition. Instead, the entire temperature interval $T\in[T^{\textrm{IBG}},T_{r}]$ should be considered as a region of instability.
@@
Since the total number of particles is conserved while the densities differ between phases, the phase transition is accompanied by discontinuities in volume and entropy, as shown in Fig. 3(b). As is well known, this behavior corresponds to a first order phase transition Huang [1987]. In particular, the entropy discontinuity leads to a nonzero latent heat. In addition, the order parameter also exhibits a discontinuous jump, as shown in Fig. 1. Employing this fact, we can clarify the volume discontinuity from a physical standpoint. Since the condensate does not contribute to the pressure, and a transition occurs at constant pressure, the loss of the latter is compensated by a decrease in volume.
@@
As we can see in Fig. 4, there is no crucial evidence (for example, $C>0$) pointing on any instability of the SSED compared to the PA. Neither the entropy nor the specific heat capacity in the normal state have linear corrections in interaction Akhiezer and Peletminskii [1981], Bulakhov et al. [2021]. Therefore, these quantities coincide with those of the ideal Bose gas.
@@
Its temperature dependence is presented in Fig. 5, where the significant qualitative differences between the two approximations become apparent. The PA exhibits a positive and generally well-behaved compressibility (with divergence only in the region of instability), as also reported in Ota and Giorgini [2020], Mordini et al. [2020], Spada et al. [2022]. In contrast, for the SSED, the compressibility is negative almost everywhere, diverging to minus infinity as $T\to 0$. Thus, the equilibrium state corresponding to the SSED is unstable.
@@
To summarize, we observe that almost all thermodynamic quantities exhibit well-behaved properties within both the HFB approach and the PA. Furthermore, the latter, which accounts solely for the single-particle condensate, predicts the positive compressibility. However, since the HFB approach does not admit a solution of (27) corresponding to a pure single-particle condensate Poluéktov [2002], Peletminskii et al. [2013], and the pair condensate (SSED) is unstable, the only possible stable state can be a mixture of single-particle and pair condensates , as also discussed below.
6 Discussion of results
@@
We have employed the energy-functional formulation to derive the self-consistency equations of the Hartree-Fock-Bogoliubov (HFB) approach. In the limit of contact interactions, these equations exhibit ultraviolet divergence, which are typically addressed by regularizing the corresponding integrals or renormalizing the coupling constant. We have obtained a new solution to the HFB self-consistency equations in which the divergence self-eliminates (the solution with self-eliminating divergence, SSED). This solution possesses several distinct features. Specifically, it predicts the existence of a condensate in the form of correlated pairs only, with zero depletion and pressure at absolute zero temperature. The vanishing pressure at zero temperature can be explained by treating the divergent term as an effective attraction between an individual pair of particles, balanced by cumulative repulsion exerted by all other particles in the gas on the given pair. We have analyzed the temperature behavior of key thermodynamic quantities in the normal state and in the degenerate states described both by the obtained SSED and Popov approximation (PA). Fig. 1 shows a good agreement of the PA and Monte-Carlo. The respective comparison with experimental data presented in A.
The common feature of both mean-field approaches is the retrograde behaviour of the condensate density, which is also observed in Ref. Prokof’ev et al. [2004] for the PA, which in turn affects other thermodynamic quantities. Nevertheless, this behaviour allows us to conclude that, due to interactions, the transition temperature shifts to higher values compared to that for an ideal gas, with the HFB approach yielding a larger shift than the PA. This positive shift of the transition temperature has also been reported in previous mean-field studies Vianello and Salasnich [2024], Arnold and Moore [2001] and Monte Carlo simulations Prokof’ev et al. [2004], Ota and Giorgini [2020], Spada et al. [2022]. In both cases, we propose to treat the phase transition between the normal and degenerate state as a first-order one. Unfortunately, the isothermic compressibility for SSED appears to be negative, which signals its instability. We emphasize that the differences among all the considered approaches, including the ideal gas case, become more pronounced as the temperature approaches $T^{\textrm{IBG}}$, which facilitates experimental verification. At the same time, it should be noted that both approaches fail to provide an adequately accurate description of thermodynamics in the vicinity of the critical temperature $T^{\textrm{IBG}}$. Therefore, an advanced possible theory that goes beyond the current mean-field approximations (PA and SSED) should not give rise to retrograde behavior and may change the nature of the phase transition.
@@
Some additional comments should also be made on certain important aspects of our study. First, although model non-contact potentials are not explicitly considered in this paper, some conclusions can nevertheless be drawn regarding their influence on the obtained results. As noted in Section 4, while they allow for a straightforward removal of divergences, the results lack stability with respect to variations in the interaction range and the choice of model potential. Furthermore, even within framework of Bogoliubov’s quadratic approximation, all solutions involving local potentials are also unstable in the same sense Bulakhov et al. [2018]. Within the HFB approach, as indicated by (34), the gap lies in the range $E_{0}\in(0,2n_{0}U]$, since PA is gapless and SSED corresponds to $E_{0}=2n_{0}U$. This allows a qualitative conclusion that for non-contact potentials, the expected results will be bounded between the curves obtained from the SSED and the PA.
@@
In connection with the Nambu-Goldstone (NG) theorem Nambu [1960], Goldstone [1961] and the Pines–Hugenholtz relation Hugenholtz and Pines [1959], it is commonly believed that the presence of a gap in the single-particle excitation spectrum is a drawback of the self-consistent HFB approach. Numerous attempts have been made to artificially remove the gap in order to satisfy this relation. In particular, the most well-known gapless PA neglects anomalous averages, and in some literature sources Spada et al. [2022], the HFB approach is even considered to be valid only in the linear order in interaction. However, as we previously pointed out, these do not represent self-consistent solutions to the coupled HFB equations. We report a unique self-consistent solution to the HFB equations for a contact interaction that resolves the divergence issue without introducing additional assumptions or artificial regularization procedures. At sufficiently high temperatures, this solution exhibits qualitative agreement with the PA, except for the isothermal compressibility.
@@
Therefore, the gathered evidence encourages us to reconsider the nature of the energy gap in the HFB approach. On the one hand, the presence of gapless (phonon) excitations is not disputed. On the other hand, the HFB approach, along with the PA artificially derived from it, represent an attempt to construct a quasiparticle description of an interacting system in terms of a single type of single-particle excitations. This holds true if the spectra of single-particle and collective excitations coincide Gavoret and Nozières [1964]. However, it has been recently demonstrated that the poles of the single-particle and two-particle Green’s functions in a Bose system with condensate do not coincide Kita [2009, 2011]. Therefore, when both types of excitations predominate in the system, the application of the HFB framework leads to a unified spectrum of single-particle excitations hybridized with collective (pair) excitations, inheriting a gap from the latter. The presence of a gap in the two-particle excitations, however, is not an artifact, as it can be associated with the energy required to break a pair.
@@
The NG theorem states that the excitation spectrum arising in a system with spontaneously broken symmetry must be gapless. This is equivalent to the statement that it should cost zero energy to excite the lowest mode (the Goldstone mode). While the direct application of the NG theorem to quasiparticles, which are not physical particles, remains debated Sütő and Szépfalusy [2008], we highlight the following observation. For the SSED, where $\Delta=0$, the second equation in (21) shows that the energy change associated with the emergence of a single correlation is zero. Therefore, when pair correlations dominate in the ground state, as described by the solution we have found, the lowest mode is collective, and its excitation costs no energy.
@@
Thus, as we have already noted, a system of interacting particles in the presence of a pair condensate requires an advanced description, which, in our view, should also reproduce a spectrum with at least two excitation branches: a phonon-like mode and a gapped mode. Similar conclusions have been drawn in Refs. Bogolubov and Bogolubov Jr [2010], Poluektov and Soroka [2023]. It should be emphasized that pure single-particle or pair condensates do not exist – only their mixture does. The existence of a purely single-particle condensate, as assumed in the PA, is ruled out by the system of HFB equations (27) (see also Poluéktov [2002], Peletminskii et al. [2013]). As for a pure pair condensate described by the SSED, it is found to be unstable based on the compressibility analysis. This result holds for the contact interaction model. In the case of finite-range interaction potentials, the corresponding HFB solutions also exhibit a gap, with $E_{0}<2n_{0}U$, and the equilibrium state likewise represents only a mixture of the two types of condensates. Such a structure of the ground state is also relevant in the context of superfluid helium, where the phonon-roton spectrum emerges from the hybridization of excitations of different physical natures and the single-particle condensate fraction is about 10% or less Sosnick et al. [1989], Bogoyavlenskii et al. [1990], Moroni and Boninsegni [2004], Glyde et al. [2011].